OpenStudy (anonymous):

help ive been stuck on this and don't get it -(5x-4)=2+8(1-2x)

OpenStudy (amistre64):

distribute the 8 and collect like terms

OpenStudy (amistre64):

might want to distribute that -1 as well

OpenStudy (amistre64):

-(5x-4) = 2 + 8(1-2x) -5x +4 = 2 + 8 -16x the rest should be pretty basic

OpenStudy (anonymous):

thank you